# Adaptive debiased SGD in high-dimensional GLMs with steaming data

### v1

@@ Abstract
Online statistical inference facilitates real-time analysis of sequentially collected data, making it different from traditional methods that rely on static datasets. This paper introduces a novel approach to online inference in high-dimensional generalized linear models, where we update regression coefficient estimates and their standard errors upon each new data arrival. In contrast to existing methods that either require full dataset access or large-dimensional summary statistics storage, our method operates in a single-pass mode, significantly reducing both time and space complexity. The core of our methodological innovation lies in an adaptive stochastic gradient descent algorithm tailored for dynamic objective functions, coupled with a novel online debiasing procedure. This allows us to maintain low-dimensional summary statistics while effectively controlling optimization errors introduced by the dynamically changing loss functions. We demonstrate, both methodologically and theoretically, that our method, termed the Approximated Debiased Lasso (ADL), not only mitigates the need for the bounded individual probability condition but also significantly improves numerical performance. Numerical experiments demonstrate that the proposed ADL method consistently exhibits robust performance across various covariance matrix structures.
@@ 1 Introduction
Online statistical inference refers to methods and algorithms used for making inference about population parameters in a real-time environment, where data is collected sequentially rather than being static as in conventional settings. More specifically, these methods update parameter estimates and quantify their uncertainty along with sequentially accrued data information. As opposed to offline learning techniques that train over a static big dataset, online learning algorithms are expected to be computationally more efficient while achieving similar statistical performances to their offline counterparts. Most existing works in the field of online statistical inference focus on low-dimensional settings where the sample size ($n$) is much larger than the number of parameters ($p$). Their main idea is to save and update a $p$ by $p$ information matrix for quantifying the uncertainty in parameter estimation (Schifano et al.,, 2016; Luo and Song,, 2020; Luo et al., 2023b, ). Nevertheless, this idea is not directly extendable to high-dimensional settings where $p\gg n$ because the space complexity of saving such a huge matrix far exceeds that of saving the raw data. It is worth noting that most existing work in online high-dimensional settings either requires the availability of an entire raw dataset or the storage of summary statistics of large dimensions (Shi et al.,, 2020; Deshpande et al.,, 2023; Luo et al., 2023a, ).
@@
In online inference, one of the main reasons to store the information matrix is that this quantity is pivotal in statistical inference and one can use such information directly to construct confidence interval for the target parameter(s). Moreover, online algorithms that leverage this second-order information can provide more accurate direction and step size towards the minimum, learning to faster convergence in some scenarios (Toulis and Airold,, 2017). However, as mentioned earlier, storing a $p\times p$ information matrix is generally not preferred in high-dimensional settings. As an alternative approach, the stochastic gradient descent (SGD) algorithm (Robbins and Monro,, 1951) fits the online setting well due to its computational and memory efficiency. In particular, SGD only requires one pass over the data, and its space complexity can reach $\mathcal{O}(p)$. The pioneering work that uses SGD for statistical inference can be traced back to Polyak and Juditsky, (1992) and it has also been applied to the low-dimensional models thereafter (Toulis and Airold,, 2017; Fang,, 2019). Recently, Chen et al., (2020) incorporate the SGD into high-dimensional inference task. Specifically, they combine the regularization annealed epoch dual averaging (RADAR) algorithm, one variant of SGD (Agarwal et al.,, 2012), with debiasing techniques (Zhang and Zhang,, 2014; van de Geer et al.,, 2014) in the linear models. However, their method requires access to the entire dataset in the debiasing step. Lately, Han et al., (2024) propose a new online debiased SGD algorithm in high-dimensional linear models that is a one-pass algorithm with space complexity $\mathcal{O}(p)$. While this algorithm has many desirable properties, it is not directly applicable to high-dimensional nonlinear models such as generalized linear models. In summary, to the best of our knowledge, existing online inference methods for high-dimensional generalized linear models are either not one-pass or require a space complexity of $\mathcal{O}(p^{2})$.
@@
Our contributions can be summarized from both methodological and theoretical aspects. Methodologically, we propose an online inference method for high-dimensional inference, named the approximated debiased lasso (ADL). The proposed method offers two main computational benefits. First, it is designed to operate in a single-pass mode, resulting in low time complexity. Second, this method does not require storing the raw data or large summary statistics, and its space complexity is only of $\mathcal{O}(p)$, which is comparable to the complexity of storing a single data point. Thus, our method can handle ultra-high dimensional data effectively. In our real data analysis, we demonstrate that our method can handle datasets with millions of covariates on a standard personal computer. By contrast, alternative methods requiring access to the entire raw dataset or large summary statistics fail in similar scenarios due to memory constraints. A detailed comparison between our methods and relevant existing literature is presented in Table 1.
@@ 2 Methodology
In the offline setting, one of the most popular approaches is the debiased lasso method (Zhang and Zhang,, 2014; van de Geer et al.,, 2014). The debiased lasso method consists of two steps: obtain the lasso estimator and correct the bias in the lasso estimator. Under the generalized linear model, we denote the negative log-likelihood function as $F_{n}(\bm{\beta})=({1}/{n})\sum_{i=1}^{n}\left\{-y_{i}(\bm{x}^{\intercal}_{i}\bm{\beta})+\Phi(\bm{x}^{\intercal}_{i}\bm{\beta})\right\}$, and specify its $\ell_{1}$-penalized form as follows:
@@
where $f_{n}(\widetilde{\bm{\beta}}^{(n)})=\nabla F_{n}(\widetilde{\bm{\beta}}^{(n)})$ is the gradient of the negative log-likelihood function, and $\widetilde{\bm{\gamma}}^{(n)}_{\cdot,j}$, known as the the nodewise lasso estimator, serves as an estimator of the $j$-th column in the inverse of Hessian matrix $\nabla^{2}F_{n}(\widetilde{\bm{\beta}}^{(n)})$. Zhang and Zhang, (2014) and van de Geer et al., (2014) show that such a debiased estimator converges to the normal distribution under the linear model and generalized linear model respectively.
@@
In the following, we develop an online method to conduct statistical inference. We first introduce the RADAR algorithm used to conduct online lasso estimation in Section 2.1. Specifically, we develop an adaptive RADAR algorithm to find online alternatives to replace $\widetilde{\bm{\gamma}}^{(n)}_{\cdot,j}$ in Section 2.2. We then propose an approximated debiasing technique to replace $f_{n}(\widetilde{\bm{\beta}}^{(n)})$. The resulting debiased estimator, as well as its standard error, are presented in Sections 2.32.4. For illustration, we present the entire procedure in Figure 1. The online lasso estimation algorithm is marked in blue. Our proposed method, including the adaptive RADAR and approximated debiasing technique, is highlighted in orange. We wish to emphasize that our procedure is one-pass and only requires $\mathcal{O}(p)$ space complexity.
@@ 2.1 Online lasso with RADAR
In this subsection, we introduce an online one-pass approach to finding the lasso estimator without retrieving or recomputing with historical data. A natural choice that serves this purpose is the SGD (Robbins and Monro,, 1951). Although SGD can obtain an estimator of $\bm{\beta}^{*}$, its convergence rate is unsatisfactory in high-dimensional settings. For example, Han et al., (2024) show that interval estimation using the vanilla SGD may not have desirable performance such as attaining nominal coverage probability. To address this issue, we adopt the regularization annealed epoch dual averaging (RADAR) algorithm (Agarwal et al.,, 2012), a variant of the vanilla SGD.
@@
It is worth noting that $\widehat{{\bm{\beta}}}^{(n)}$, the final output from the RADAR algorithm, is no longer the minimizer of $F_{n}(\bm{\beta};\lambda_{n})$ defined in (2). In fact, without accessing the entire dataset, it is hard to achieve the optimality as the offline estimator. Recall $\widetilde{\bm{\beta}}^{(n)}$ is the minimizer of $F_{n}(\bm{\beta};\lambda_{n})$. Instead of the statistical error $\lVert\widetilde{\bm{\beta}}^{(n)}-\bm{\beta}^{*}\lVert_{1}$, we need to consider the optimization error from RADAR, that is $\lVert\widehat{{\bm{\beta}}}^{(n)}-\bm{\beta}^{*}\lVert_{1}$. Under some regularity conditions, we can conclude that the optimization error $\lVert\widehat{{\bm{\beta}}}^{(n)}-\bm{\beta}^{*}\lVert_{1}$ achieves the same rate as the statistical error $\lVert\widetilde{\bm{\beta}}^{(n)}-\bm{\beta}^{*}\lVert_{1}$. A detailed description of the RADAR algorithm is included in the supplementary material.
@@ 2.2 Online nodewise lasso with adaptive RADAR
${\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}={\widehat{\bm{\gamma}}_{\cdot,j}}^{(n_{k}^{\prime})},\ \text{if}\ n_{k}^{\prime}\leq i<n_{k+1}^{\prime},$ (7)
@@
The adaptive RADAR is more general than the original RADAR algorithm. Specifically, let $\mathcal{F}_{k}^{\prime}=\sigma\left(\{(\bm{x}_{i},y_{i})\}_{i=1}^{n_{k}^{\prime}}\right)$, that is the $\sigma$-field generated from $\{(\bm{x}_{i},y_{i})\}_{i=1}^{n_{k}^{\prime}}$, and $\bar{G}_{k}(\bm{r})=\mathbb{E}\left(\ddot{\Phi}(\bm{X}^{\intercal}\widehat{\bm{\beta}}^{(n_{k-1}^{\prime})})(\bm{X}^{\intercal}\bm{r})^{2}\big{\lvert}\mathcal{F}_{k-1}^{\prime}\right)$. In the $k$-th epoch, the stochastic gradient defined in (6) comes from the objective function
@@ 2.3 Approximated debiasing techniques
Note that the Hessian matrices $\bm{x}_{i}\ddot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})\bm{x}^{\intercal}_{i},\ i\in[n]$ are $p\times p$ matrices, which requires a large amount of memory capacity. To ensure the superiority of our method in terms of the space complexity, we construct the following summary statistics. Letting $l=\min\{L:n_{L}\geq n_{1}^{\prime}\}$, we compute
$\bm{a}_{1}^{(k)}=\sum_{i=n_{l}+1}^{k}\bm{x}_{i}\left\{-y_{i}+\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})\right\},\ \bm{a}_{2}^{(k)}=\sum_{i=n_{l}+1}^{k}\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}\ddot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})\bm{x}_{i},$
$a_{3}^{(k)}=\sum_{i=n_{l}+1}^{k}\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}\ddot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)},\ {a}_{4}^{(k)}=\sum_{i=n_{l}+1}^{k}\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}(\bm{x}_{i})_{j}\left\{\ddot{\Phi}(\bm{x}_{i}^{\intercal}\widehat{\bm{\beta}}^{(i)})\right\},$ (12)
where $\bm{a}_{1}^{(k)}$ and $\bm{a}_{2}^{(k)}$ are $p$-dimensional vectors while the other elements $a_{3}^{(k)}$ and $a_{4}^{(k)}$ are scalars. Note that the above summary statistics are computed from $i=n_{l}+1$ rather than $i=1$. The main purpose is to guarantee the existence of ${\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}$ and $\widehat{\bm{\beta}}^{(i)}$.
Based on the statistics defined in (12), we define the approximated debiased lasso (ADL) estimator as
${\widehat{\bm{\beta}}}^{(k)}_{j,\text{de}}:=\widehat{{\bm{\beta}}}^{(k)}_{j}-\frac{\{\bm{a}^{(k)}_{1}\}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(k)}+\{\bm{a}^{(k)}_{2}\}^{\intercal}\widehat{{\bm{\beta}}}^{(k)}-a_{3}^{(k)}}{{a}_{4}^{(k)}},\ k>n_{l}.$ (13)
One key advantage of (13) is that we can update these summary statistics recursively instead of storing the raw data in memory. For example, when the $k$-th data arrives, we compute
$\bm{a}_{1}^{(k)}=\bm{a}_{1}^{(k-1)}+\bm{x}_{k}\left\{-y_{k}+\dot{\Phi}(\bm{x}^{\intercal}_{k}\widehat{\bm{\beta}}^{(k)})\right\}.$ (14)
Then, following the update as in (14), the calculation of ${\widehat{\bm{\beta}}}^{(k)}_{j,\text{de}}$ in (13) is one-pass with $\mathcal{O}(p)$ space complexity. For ease of illustration, a diagram that depicts the alignment of the three key steps in online estimation and inference procedure is presented in Figure 2.
Figure 2: The alignment of three key steps in constructing the ADL estimator.

### v2

# Adaptive Debiased SGD in High-dimensional Generalized Linear Models with Streaming Data

@@ Abstract
Online statistical inference facilitates real-time analysis of sequentially collected data, making it different from traditional methods that rely on static datasets. This paper introduces a novel approach to online inference in high-dimensional generalized linear models, where we update regression coefficient estimates and their standard errors upon each new data arrival. In contrast to existing methods that either require full dataset access or large-dimensional summary statistics storage, our method operates in a single-pass mode, significantly reducing both time and space complexity. The core of our methodological innovation lies in an adaptive stochastic gradient descent algorithm tailored for dynamic objective functions, coupled with a novel online debiasing procedure. This allows us to maintain low-dimensional summary statistics while effectively controlling optimization errors introduced by the dynamically changing loss functions. We demonstrate, both methodologically and theoretically, that our method, termed the Adaptive Debiased Lasso (ADL), not only mitigates the need for the bounded individual probability condition but also significantly improves numerical performance. Numerical experiments demonstrate that the proposed ADL method consistently exhibits robust performance across various covariance matrix structures.
@@ 1 Introduction
Online statistical inference refers to methods and algorithms used for making inference about population parameters in a real-time environment, where data is collected sequentially rather than being static as in conventional settings. More specifically, these methods update parameter estimates and quantify their uncertainty along with sequentially accrued data information. As opposed to offline learning techniques that train over a static big dataset, online learning algorithms are expected to be computationally more efficient while achieving similar statistical performances to their offline counterparts. Most existing works in the field of online statistical inference focus on low-dimensional settings where the sample size ($n$) is much larger than the number of parameters ($p$). Their main idea is to save and update a $p$ by $p$ information matrix for quantifying the uncertainty in parameter estimation (Schifano et al.,, 2016; Luo and Song,, 2020; Luo et al., 2023b, ). Nevertheless, this idea is not directly extendable to high-dimensional settings where $p\gg n$ because the space complexity of saving such a huge matrix far exceeds that of saving the raw data. It is worth noting that most existing work in online high-dimensional settings either requires the availability of an entire raw dataset or the storage of summary statistics of large dimensions (Shi et al.,, 2021; Deshpande et al.,, 2023; Luo et al., 2023a, ).
@@
In online inference, one of the main reasons to store the information matrix is that this quantity is pivotal in statistical inference and one can use such information directly to construct confidence interval for the target parameter(s). Moreover, online algorithms that leverage this second-order information can provide more accurate direction and step size towards the minimum, learning to faster convergence in some scenarios (Toulis and Airold,, 2017). However, as mentioned earlier, storing a $p\times p$ information matrix is generally not preferred in high-dimensional settings. As an alternative approach, the stochastic gradient descent (SGD) algorithm (Robbins and Monro,, 1951) fits the online setting well due to its computational and memory efficiency. In particular, SGD only requires one pass over the data, and its space complexity can reach $\mathcal{O}(p)$. The pioneering work that uses SGD for statistical inference can be traced back to Polyak and Juditsky, (1992) and it has also been applied to the low-dimensional models thereafter (Toulis and Airold,, 2017; Fang,, 2019; Chen et al.,, 2024). Recently, Chen et al., (2020) incorporate the SGD into high-dimensional inference task. Specifically, they combine the regularization annealed epoch dual averaging (RADAR) algorithm, one variant of SGD (Agarwal et al.,, 2012), with debiasing techniques (Zhang and Zhang,, 2014; van de Geer et al.,, 2014) in the linear models. However, their method requires access to the entire dataset in the debiasing step. Lately, Han et al., (2024) propose a new online debiased SGD algorithm in high-dimensional linear models that is a one-pass algorithm with space complexity $\mathcal{O}(p)$. While this algorithm has many desirable properties, it is not directly applicable to high-dimensional nonlinear models such as generalized linear models. In summary, to the best of our knowledge, existing online inference methods for high-dimensional generalized linear models are either not one-pass or require a space complexity of $\mathcal{O}(p^{2})$.
@@
Our contributions can be summarized from both methodological and theoretical aspects. Methodologically, we propose an online inference method for high-dimensional inference, named the adaptive debiased lasso (ADL). The proposed method offers two main computational benefits. First, it is designed to operate in a single-pass mode, resulting in low time complexity. Second, this method does not require storing the raw data or large summary statistics, and its space complexity is only of $\mathcal{O}(p)$, which is comparable to the complexity of storing a single data point. Thus, our method can handle ultra-high dimensional data effectively. In our real data analysis, we demonstrate that our method can handle datasets with millions of covariates on a standard personal computer. By contrast, alternative methods requiring access to the entire raw dataset or large summary statistics fail in similar scenarios due to memory constraints. A detailed comparison between our method and relevant existing literature is presented in Table 1.
@@ 2 Methodology
In the offline setting, one of the most popular approaches is the debiased lasso method (Zhang and Zhang,, 2014; van de Geer et al.,, 2014). The debiased lasso method consists of two steps: obtain the lasso estimator and correct the bias in the lasso estimator. Under the generalized linear models, we denote the negative log-likelihood function as $F_{n}(\bm{\beta})=({1}/{n})\sum_{i=1}^{n}\left\{-y_{i}(\bm{x}^{\intercal}_{i}\bm{\beta})+\Phi(\bm{x}^{\intercal}_{i}\bm{\beta})\right\}$, and specify its $\ell_{1}$-penalized form as follows:
@@
where $f_{n}(\widetilde{\bm{\beta}}^{(n)})=\nabla F_{n}(\widetilde{\bm{\beta}}^{(n)})$ is the gradient of the negative log-likelihood function, and $\widetilde{\bm{\gamma}}^{(n)}_{\cdot,j}$, known as the the nodewise lasso estimator, serves as an estimator of the $j$-th column in the inverse of Hessian matrix $\nabla^{2}F_{n}(\widetilde{\bm{\beta}}^{(n)})$. Zhang and Zhang, (2014) and van de Geer et al., (2014) show that such a debiased estimator converges to the normal distribution under the linear models and generalized linear models respectively.
@@
In the following, we develop an online method to conduct statistical inference. We first introduce the RADAR algorithm used to conduct online lasso estimation in Section 2.1. Specifically, we develop an adaptive RADAR algorithm to find online alternatives to replace $\widetilde{\bm{\gamma}}^{(n)}_{\cdot,j}$ in Section 2.2. We then propose an approximated debiasing technique to replace $f_{n}(\widetilde{\bm{\beta}}^{(n)})$. The resulting estimator, named adaptive debiased lasso (ADL) estimator, as well as its standard error, are presented in Sections 2.32.4. For illustration, we present the entire procedure in Figure 1. The online lasso estimation algorithm is marked in blue. Our proposed method, including the adaptive RADAR and approximated debiasing technique, is highlighted in orange. We wish to emphasize that our procedure is one-pass and only requires $\mathcal{O}(p)$ space complexity.
@@ 2.1 Online lasso with RADAR
In this subsection, we introduce an online one-pass approach to finding the lasso estimator without retrieving or recomputing historical data. A natural choice that serves this purpose is the SGD (Robbins and Monro,, 1951). Although SGD can obtain an estimator of $\bm{\beta}^{*}$, its convergence rate is unsatisfactory in high-dimensional settings. For example, Han et al., (2024) show that interval estimation using the vanilla SGD may not have desirable performance such as attaining nominal coverage probability. To address this issue, we adopt the regularization annealed epoch dual averaging (RADAR) algorithm (Agarwal et al.,, 2012), a variant of the vanilla SGD.
@@
It is worth noting that $\widehat{{\bm{\beta}}}^{(n)}$, the final output from the RADAR algorithm, is no longer the minimizer of $F_{n}(\bm{\beta};\lambda_{n})$ defined in (2). In fact, without accessing the entire dataset, it is hard to achieve the optimality as the offline estimator. Recall $\widetilde{\bm{\beta}}^{(n)}$ is the minimizer of $F_{n}(\bm{\beta};\lambda_{n})$. Instead of the statistical error $\lVert\widetilde{\bm{\beta}}^{(n)}-\bm{\beta}^{*}\lVert_{1}$, we need to consider the optimization error from the RADAR algorithm, that is $\lVert\widehat{{\bm{\beta}}}^{(n)}-\bm{\beta}^{*}\lVert_{1}$. Under some regularity conditions, we can conclude that the optimization error $\lVert\widehat{{\bm{\beta}}}^{(n)}-\bm{\beta}^{*}\lVert_{1}$ achieves the same rate as the statistical error $\lVert\widetilde{\bm{\beta}}^{(n)}-\bm{\beta}^{*}\lVert_{1}$. A detailed description of the RADAR algorithm is included in the supplementary material.
@@ 2.2 Online nodewise lasso with adaptive RADAR
${\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}={\widehat{\bm{\gamma}}_{\cdot,j}}^{(n_{k}^{\prime})},\ \text{if}\ n_{k}^{\prime}+1\leq i<n_{k+1}^{\prime},$ (7)
@@
The adaptive RADAR algorithm is more general than the original RADAR algorithm. Specifically, let $\mathcal{F}_{k}^{\prime}=\sigma\left(\{(\bm{x}_{i},y_{i})\}_{i=1}^{n_{k}^{\prime}}\right)$, that is the $\sigma$-field generated from $\{(\bm{x}_{i},y_{i})\}_{i=1}^{n_{k}^{\prime}}$, and $\bar{G}_{k}(\bm{r})=\mathbb{E}\left(\ddot{\Phi}(\bm{X}^{\intercal}\widehat{\bm{\beta}}^{(n_{k-1}^{\prime})})(\bm{X}^{\intercal}\bm{r})^{2}\big{\lvert}\mathcal{F}_{k-1}^{\prime}\right)$. In the $k$-th epoch, the stochastic gradient defined in (6) comes from the objective function
@@ 2.3 Approximated debiasing techniques
Note that the Hessian matrices $\bm{x}_{i}\ddot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})\bm{x}^{\intercal}_{i},\ i\in[n]$ are $p\times p$ matrices, which require a large amount of memory capacity. To ensure the superiority of our method in terms of the space complexity, we construct the following summary statistics. Letting $l=\min\{{L}:n_{L}\geq n_{1}^{\prime}\}$, we compute
$\bm{a}_{1}^{(m)}=\sum_{i=n_{l}+1}^{m}\bm{x}_{i}\left\{-y_{i}+\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})\right\},\ \bm{a}_{2}^{(m)}=\sum_{i=n_{l}+1}^{m}\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}\ddot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})\bm{x}_{i},$
$a_{3}^{(m)}=\sum_{i=n_{l}+1}^{m}\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}\ddot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)},\ {a}_{4}^{(m)}=\sum_{i=n_{l}+1}^{m}\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}(\bm{x}_{i})_{j}\left\{\ddot{\Phi}(\bm{x}_{i}^{\intercal}\widehat{\bm{\beta}}^{(i)})\right\},$ (12)
where $\bm{a}_{1}^{(m)}$ and $\bm{a}_{2}^{(m)}$ are $p$-dimensional vectors while the other elements $a_{3}^{(m)}$ and $a_{4}^{(m)}$ are scalars. The above summary statistics are computed from $i=n_{l}+1$ rather than $i=1$. The main purpose is to guarantee the existence of ${\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}$ and $\widehat{\bm{\beta}}^{(i)}$.
Based on the statistics defined in (12), we define the adaptive debiased lasso (ADL) estimator as
${\widehat{\bm{\beta}}}^{(m)}_{j,\text{de}}:=\widehat{{\bm{\beta}}}^{(m)}_{j}-\frac{\{\bm{a}^{(k)}_{1}\}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(m)}+\{\bm{a}^{(m)}_{2}\}^{\intercal}\widehat{{\bm{\beta}}}^{(k)}-a_{3}^{(m)}}{{a}_{4}^{(m)}},\ m>n_{l}.$ (13)
One key advantage of (13) is that we can update these summary statistics recursively instead of storing the raw data in memory. For example, when the $m$-th data arrives, we compute
$\bm{a}_{1}^{(m)}=\bm{a}_{1}^{(m-1)}+\bm{x}_{m}\left\{-y_{m}+\dot{\Phi}(\bm{x}^{\intercal}_{m}\widehat{\bm{\beta}}^{(m)})\right\}.$ (14)
Then, following the update as in (14), the calculation of ${\widehat{\bm{\beta}}}^{(m)}_{j,\text{de}}$ in (13) is one-pass with $\mathcal{O}(p)$ space complexity. For ease of illustration, a diagram that depicts the alignment of the three key steps in online estimation and inference procedure is presented in Figure 2.
Figure 2: The timeline of three key steps in constructing the ADL estimator.